OpenStudy (anonymous):

\\HELP// Write an equation in slope-intercept form of the line that passes through (1,-3) and is parallel to the graph of y+2=4(x-1)

OpenStudy (anonymous):

This is actually very simple. Because the line is parallel, the slope is the same. To find the slope of y + 2 = 4(x - 1), convert the function into slope intercept form, that is (y = mx + b) So, the equation becomes y + 2 = 4x - 4 y = 4x -4 - 2 y = 4x - 6 Slope is 4. Then substitute the points (1, -3) with the slope of 4 to find the b in y = mx + b -3 = (4)(1) + b -3 = 4 + b -3 + 4 = b 1 = b The final question is y = 4x + 1

OpenStudy (anonymous):

Sorry, I meant to say "The final EQUATION is y = 4x + 1 Also, keep in mind that linear equations (y = mx + b or y2 - y1 = m(x2 - x1) that are parallel only differ in b. The same is true for the above question. The b part in y = mx + b determines how much up or down the line is in the graph. For a line to be parallel, it just needs to be a little more up or down.

OpenStudy (anonymous):

thanks it really helped :)